# Circle Squaring with Pieces of Small Boundary and Low Borel Complexity

### v3

@@ 6.2. Rounding a Sequence of Flows on a Toast Sequence
Let $\phi$ and $\psi$ be flows in $G_{d}$, $M$ be a non-negative real number and $D$ be a subset of $\mathbb{T}^{k}$ such that $|\partial_{E}S|\leqslant M$ for every component $S$ of $G_{d}\upharpoonright D$ and any two components of $G_{d}\upharpoonright D$ are at distance at least three in $G_{d}$. Then $\|{{\mathcal{E}}}^{\phi,\psi}_{D}\|_{\infty}\leqslant M\|\phi-\psi\|_{\infty}$.
@@
For each $i\geqslant 1$, let $m_{i}$ and $m_{i}^{\prime}$ be the minimal integers such that
@@
Let $f=f_{J\widehat{K}\widehat{L}}$ be the rounding of the sequence $(f_{m_{i}},f_{m_{i}},f_{m_{i}^{\prime}})_{i=1}^{\infty}$ of flows on the toast sequence $(J_{i},\widehat{K}_{i},\widehat{L}_{i})_{i=1}^{\infty}$. (Recall that the sets $\widehat{K}_{i}$ and $\widehat{L}_{i}$ were defined before Lemma 5.15 and $(J_{i},\widehat{K}_{i},\widehat{L}_{i})_{i=1}^{\infty}$ is a toast sequence by the lemma; also, $f_{m}$ for $m\in{\mathbbm{N}}$ is the flow returned by Lemma 3.1.) We will use this flow $f$ to obtain the desired equidecomposition between $A$ and $B$.
@@ 7.1. Dimension of Boundaries
Here we prove Part 1.1 of Theorem 1.3 that the upper Minkowski dimension of the obtained pieces is strictly smaller than $k$. Specifically, we show that, for every $\vec{\gamma}\in\{-1,0,1\}^{d}$ and $|\ell|\leqslant\|f\|_{\infty}$, the upper Minkowski dimension of the boundary of $Z_{\vec{\gamma},\ell}^{f}$ is at most $k-\zeta$, where
@@
Let $C_{1}$ be a constant satisfying Lemma 7.4 with $t=1$ for at least half (in measure) of the choices of $(\boldsymbol{x}_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,\boldsymbol{x}_{d})\in({\mathbbm{T}}^{k})^{d}$. One way to see its existence (without checking whether the proof of Lemma 7.4 gives some effective bounds on $C_{1}$) is as follows. For every real $C$, the set ${\mathcal{X}}_{C}$ of sequences $(\boldsymbol{x}_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,\boldsymbol{x}_{d})\in({\mathbbm{T}}^{k})^{d}$ for which Lemma 7.4 holds for this $C$ (with $t=1$) is an analytic subset, since it is the projection of the Borel subset of $({\mathbbm{T}}^{k})^{d}\times\Omega$ of points satisfying (7.12), where $\Omega\subseteq{\mathbbm{T}}^{k}\times{\mathbbm{N}}\times{\mathbbm{T}}^{2k}$ is the subspace encoding all suitable triples $(\boldsymbol{u},r,I)$ equipped with the standard (Polish) topology. As every analytic subset is universally measurable (see e.g. [Kechris:cdst]*Theorem 21.10) and the countable nested union $\bigcup_{i=1}^{\infty}{\mathcal{X}}_{i}\subseteq{\mathbbm{T}}^{k}$ has Lebesgue measure 1, by the countable additivity there is $i\in{\mathbbm{N}}$ such that the measure of ${\mathcal{X}}_{i}$ is at least $1/2$ and we can take $C_{1}$ to be equal to this $i$.
We will use the following estimates that hold by choosing the constants $c_{1}>0$ and $C_{2}\gg C_{1}$ suitably:
@@
(7.15) $(2R_{i}+1)^{d}\gamma_{i}>C_{1}\log^{k+d+1}(2R_{i}),\quad\mbox{for every integer $i\geqslant 1$.}$
Let us briefly check that suitable $c_{1}$ and $C_{2}$ exist. The $i$-th summand in the left-hand side of (7.14) is at most $(3r_{i})^{d}\cdot c_{1}r_{i}^{-d}\log^{-2}(r_{i})\leqslant O(c_{1}2^{-2i})$. This is summable, so (7.14) can be satisfied by choosing $c_{1}$ sufficiently small. Also, the left-hand size of (7.15) is at least $R_{i}^{d}\gamma_{i}=C_{2}^{d}\cdot c_{1}\cdot\log^{d+r+1}(r_{i})$. As the exponent at the logarithm matches that in the right-hand side of (7.15), this constraint can be satisfied by taking $C_{2}$ sufficiently large.
@@
Thus the probability that $Q_{i}$ is not $r_{i}$-discrete for some $i\geqslant 1$ is at most $\sum_{i=1}^{\infty}4(2r_{i}+1)^{d}\gamma_{i}$ which is at most $1/3$ by (7.14).
@@
(7.16) $\max\big{\{}\lambda\left(\left\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial A)<\delta\right\}\right),\,\lambda\left(\left\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial B)<\delta\right\}\right)\big{\}}<\delta^{\epsilon},$
@@
By Lemma 7.7, any set of the form $Z\cap I_{i}$ for $Z\in\mathcal{Z}$ can be written as a Boolean combination of at most $r_{i}^{d\,\max\{d/\varepsilon,d+1\}+o(1)}$ translates of $A$, $B$ and $Q_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,Q_{i}$. The union of all boxes in $\mathcal{B}$ lies within $\operatorname{dist}_{\infty}$-distance at most $\delta$ from the boundary of at least one of these translates. Thus, using the volume bound, the estimate in (7.16) and that, trivially, the measure of $\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial Q_{i})\leqslant\delta\}$ is at most $4\delta$, we get by Lemma 7.7 that
$\delta^{k}|\mathcal{B}|$ $\leqslant$ $2^{k}\lambda\left(\bigcup_{Z\in\mathcal{Z}}\left\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial Z)\leqslant\delta\right\}\right)$
@@
$\delta^{k}|\mathcal{B}|\leqslant(\delta^{-4\zeta})^{d\,\max\{d/\varepsilon,\,d+1\}+\epsilon-o(1)}\leqslant\delta^{\zeta-o(1)}.$
@@ 7.2. Borel Complexity
[…]
@@
By construction (i.e. by Conclusion (7.7) of Lemma 7.2), the flows $f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}$ and $f_{J}$ coincide on every pair intersecting $\bigcup_{i=1}^{\infty}J_{i}$. Also, if an edge $\boldsymbol{u}\boldsymbol{v}$ intersects $K_{j}^{\vec{p}}$ then, by (5.17), the value of $f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}$ on this pair equals the value of $f_{JK^{\vec{p}}}$ unless there exists $i>j$ such that $\boldsymbol{u}\boldsymbol{v}$ intersects $J_{i}$ and the values of $f_{J}$ and $f_{JK^{\vec{p}}}$ on this pair differ.
Informally speaking, the value under $f_{JK^{\vec{p}}}$ is retained by the final flow, unless it is overwritten (to a different value or the same one) due to a later set $J_{i}$. Similarly, if an edge intersects $L_{j}^{\vec{p}}$, then the value of $f_{K^{\vec{p}}L^{\vec{p}}}$ on this edge is retained by $f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}$, unless it is overwritten due to $K_{i}^{\vec{p}}$ with $i>j$ intersecting this pair. Thus we have three types of flow values: $J$-values, $K$-values not overwritten by $J$, and $L$-values not overwritten by $K$.
Formally, we can express the above partition as
@@
Note that each union is over a finite set (since the involved flows are integer-valued and uniformly bounded). This proves (7.21) and completes the proof of the lemma.
@@ Appendix A Laczkovich’s Discrepancy Bound
We apply Lemma 7.4 to the first sum on the right side of (A.2) and obtain that
@@
So, if we set $\delta\leqslant(r+1)^{-d/k}$, then we get

### v4

@@ 6.2. Rounding a Sequence of Flows on a Toast Sequence
Let $\phi$ and $\psi$ be flows in $G_{d}$, $M$ be a non-negative integer and $D$ be a subset of $\mathbb{T}^{k}$ such that $|\partial_{E}S|\leqslant M$ for every component $S$ of $G_{d}\upharpoonright D$ and any two components of $G_{d}\upharpoonright D$ are at distance at least three in $G_{d}$. Then $\|{{\mathcal{E}}}^{\phi,\psi}_{D}\|_{\infty}\leqslant M\|\phi-\psi\|_{\infty}$.
@@
For each $i\geqslant 1$, let $m_{i}$ and $m_{i}^{\prime}$ be the minimal non-negative integers such that
@@
Let $f=f_{J\widehat{K}\widehat{L}}$ be the rounding of the sequence $(f_{m_{i}},f_{m_{i}},f_{m_{i}^{\prime}})_{i=1}^{\infty}$ of flows on the toast sequence $(J_{i},\widehat{K}_{i},\widehat{L}_{i})_{i=1}^{\infty}$. (Recall that the sets $\widehat{K}_{i}$ and $\widehat{L}_{i}$ were defined before Lemma 5.15 and $(J_{i},\widehat{K}_{i},\widehat{L}_{i})_{i=1}^{\infty}$ is a toast sequence by the lemma; also, $f_{m}$ for $m\geqslant 0$ is the flow returned by Lemma 3.1.) We will use this flow $f$ to obtain the desired equidecomposition between $A$ and $B$.
@@ 7.1. Dimension of Boundaries
Here we prove Part 1.1 of Theorem 1.3 that the upper Minkowski dimension of the boundaries of the obtained pieces is strictly smaller than $k$. Specifically, we show that, for every $\vec{\gamma}\in\{-1,0,1\}^{d}$ and $|\ell|\leqslant\|f\|_{\infty}$, the upper Minkowski dimension of the boundary of $Z_{\vec{\gamma},\ell}^{f}$ is at most $k-\zeta$, where
@@
Let $C_{1}$ be a constant satisfying Lemma 7.4 with $t=1$ for at least half (in measure) of the choices of $(\boldsymbol{x}_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,\boldsymbol{x}_{d})\in({\mathbbm{T}}^{k})^{d}$. One way to see its existence (without checking whether the proof of Lemma 7.4 gives some effective bounds on $C_{1}$) is as follows. For every real $C$, the set ${\mathcal{X}}_{C}$ of sequences $(\boldsymbol{x}_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,\boldsymbol{x}_{d})\in({\mathbbm{T}}^{k})^{d}$ for which Lemma 7.4 holds for this $C$ (with $t=1$) is an analytic subset, since it is the projection of the Borel subset of $({\mathbbm{T}}^{k})^{d}\times\Omega$ of points satisfying (7.12), where $\Omega\subseteq{\mathbbm{T}}^{k}\times\{2,3,\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt\}\times{\mathbbm{T}}^{2k}$ is the subspace encoding all suitable triples $(\boldsymbol{u},r,I)$ equipped with the standard (Polish) topology. As every analytic subset is universally measurable (see e.g. [Kechris:cdst]*Theorem 21.10) and the countable nested union $\bigcup_{i=1}^{\infty}{\mathcal{X}}_{i}\subseteq{\mathbbm{T}}^{k}$ has Lebesgue measure 1, by the countable additivity there is an index $i$ such that the measure of ${\mathcal{X}}_{i}$ is at least $1/2$ and we can take $C_{1}$ to be equal to this $i$.
We will use the following estimates that hold by choosing the constants $c_{1}>0$ and $C_{2}\gg C_{1}$ suitably: for every integer $i\geqslant 1$,
@@
(7.15) $(2R_{i}+1)^{d}\gamma_{i}>C_{1}\log^{k+d+1}(2R_{i}).$
Let us briefly check that suitable $c_{1}$ and $C_{2}$ exist. The $i$-th summand in the left-hand side of (7.14) is at most $(3r_{i})^{d}\cdot c_{1}r_{i}^{-d}\log^{-2}(r_{i})\leqslant O(c_{1}2^{-2i})$. This is summable, so (7.14) can be satisfied by choosing $c_{1}$ sufficiently small. Also, the left-hand size of (7.15) is at least $R_{i}^{d}\gamma_{i}=C_{2}^{d}\cdot c_{1}\cdot\log^{k+d+1}(r_{i})$. As the exponent at the logarithm matches that in the right-hand side of (7.15), this constraint can be satisfied by taking $C_{2}$ sufficiently large.
@@
Thus the probability that the set $Q_{i}$ is not $r_{i}$-discrete for some $i\geqslant 1$ is at most $\sum_{i=1}^{\infty}4(2r_{i}+1)^{d}\gamma_{i}$ which is at most $1/3$ by (7.14).
@@
(7.16) $\max\big\{\lambda\left(\left\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial A)\leqslant\delta\right\}\right),\,\lambda\left(\left\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial B)\leqslant\delta\right\}\right)\big\}<\delta^{\epsilon},$
@@
By Lemma 7.7, any set of the form $Z\cap I_{i}$ for $Z\in\mathcal{Z}$ can be written as a Boolean combination of at most $r_{i}^{d\,\max\{d/\varepsilon,d+1\}+o(1)}$ translates of $A$, $B$ and $Q_{1},\hskip 0.9pt.\hskip 0.3pt.\hskip 0.3pt.\hskip 1.5pt,Q_{i}$. The union of all boxes in $\mathcal{B}$ lies within $\operatorname{dist}_{\infty}$-distance at most $\delta$ from the boundary of at least one of these translates. Thus, using the volume bound, the estimate in (7.16) and that, trivially, the measure of $\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial Q_{i})\leqslant\delta\}$ is at most $4\delta$, we get that
$\delta^{k}|\mathcal{B}|$ $\leqslant$ $2^{k}\lambda\left(\bigcup_{Z\in\mathcal{Z}}\left\{\boldsymbol{x}:\operatorname{dist}_{\infty}(\boldsymbol{x},\partial(Z\cap I_{i}))\leqslant\delta\right\}\right)$
@@
$\delta^{k}|\mathcal{B}|\leqslant(\delta^{-4\zeta})^{d\,\max\{d/\varepsilon,\,d+1\}+o(1)}\cdot\delta^{\epsilon-o(1)}\leqslant\delta^{\zeta-o(1)}.$
@@ 7.2. Borel Complexity
Recall that each of $(J_{i},K_{i}^{\vec{p}})_{i=1}^{\infty}$ and $(K_{i}^{\vec{p}},L_{i}^{\vec{p}})_{i=1}^{\infty}$ is a toast sequence by Lemma 5.9. Also, Lemma 7.2 (with the used toast sequence $(J_{i}^{\prime},K_{i}^{\prime},L_{i}^{\prime})_{i=1}^{\infty}$ being respectively $(J_{i},K_{i}^{\vec{p}},\emptyset)_{i=1}^{\infty}$ and $(\emptyset,K_{i}^{\vec{p}},L_{i}^{\vec{p}})_{i=1}^{\infty}$) applies to each of these two new flows.
@@
By construction (i.e. by Conclusion (7.7) of Lemma 7.2), the flows $f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}$ and $f_{J}$ coincide on every pair intersecting $\bigcup_{i=1}^{\infty}J_{i}$. Also, if an edge $\boldsymbol{u}\boldsymbol{v}$ intersects $K_{j}^{\vec{p}}$ then, by (5.17), the value of $f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}$ on this edge equals the value of $f_{JK^{\vec{p}}}$ unless there exists $i>j$ such that $\boldsymbol{u}\boldsymbol{v}$ intersects $J_{i}$ when we use the value of $f_{J}$ (which may happen to coincide with the value of $f_{JK^{\vec{p}}}$ on $\boldsymbol{u}\boldsymbol{v}$). Furthermore, we can drop the restriction that $i>j$ here, since our rounding procedures for any $J_{i}$ and $J_{i^{\prime}}$ produce the same value on each edge intersecting $J_{i}\cap J_{i^{\prime}}$. Similarly, if an edge intersects $L_{j}^{\vec{p}}$, then the value of $f_{K^{\vec{p}}L^{\vec{p}}}$ on this edge is retained by $f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}$, unless it is overwritten (to a different value or the same one) due to $J_{i}$ or $K_{i}^{\vec{p}}$ with $i>j$ intersecting this edge; again we can drop the restriction that $i>j$ here.
Thus, informally speaking, our flow values are partitioned into the following three types: $J$-values, $K$-values not overwritten by $J$, and $L$-values not overwritten by $J$ nor $K$, that is, the order of precedence is $J,K,L$. Formally, we can express the above partition for every non-zero $\ell$ as
@@
Note that each union is over a finite set (since the involved flows are integer-valued and uniformly bounded), thus satisfying (7.21). The case $\ell=0$ is somewhat special since the union $\bigcup_{i=1}^{\infty}(J_{i}\cup K^{\vec{p}}_{i}\cup L^{\vec{p}}_{i})$ need not cover the whole torus. However, this case also satisfies (7.21) since $Z_{\vec{n},0}^{f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}}$ is the complement of the union of the sets $Z_{\vec{n},\ell}^{f_{J\widetilde{K}^{\vec{p}}\widetilde{L}^{\vec{p}}}}$ over all possible (finitely many) non-zero flow values $\ell$. This completes the proof of the lemma.
@@ Acknowledgements
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix A Laczkovich’s Discrepancy Bound
We apply Lemma 7.4 to the first sum on the right side of (A.2), say with $t:=1$, and obtain that
@@
So, if we set $\delta:=(r+1)^{-d/k}$ for $r\to\infty$, then we get